derivation of second order reaction 
with graph and half life of second order reaction

Grade:12th pass

1 Answers

Vikas TU
7 years ago
From the definition,
d[A]/dt = -k[A]^2
where [A] is the initial conc.
Now integerating both sides with suitably measures,
d(A)/(A)^2 = -kdt
-[(1/At – 1/A0)] = -kt
or
[(1/At – 1/A0)] = kt + c
 
For thalf,
t1/2 = 1/KA0